OpenStudy (anonymous):

anyone know what to do logarithms with a negative exponent on the TI-84 ?

OpenStudy (campbell_st):

can you give an example

OpenStudy (campbell_st):

perhaps putting the negative exponent in brackets may help

OpenStudy (anonymous):

no I'm using an old schaum's outline series for College Algebra. It's so old it still uses the tables with mantissas and characteristic. I'm trying to do the same problem with my calculator

OpenStudy (campbell_st):

an an example \[\log(5^{(-4)}) = \]

OpenStudy (campbell_st):

and you'd get -2.796

OpenStudy (anonymous):

Their problem is Find N if if log N = 7.8657 - 10 ( the minus 10 makes it negative. My ti-84 says answer is .000136623, the answer is supposed to be .00734

OpenStudy (anonymous):

sorry .00013623

OpenStudy (anonymous):

Minus 10 make it -3.8657

OpenStudy (anonymous):

I took 10^(-3.8657)

OpenStudy (campbell_st):

so you are working in base 10 logs.... so if you raise each side of the equation as a power of 10 you will get \[10^{logN} = 10^{7.8657 - 10}\] which becomes \[N = 10^{(7.8657-10)}\] so on your calculator enter 10^(7.8657 - 10) then press equals

OpenStudy (campbell_st):

and you'll find the answer matches the text

OpenStudy (anonymous):

I did get same answer doing it your way and it matches the book, but why doesn't it work in ti when I actually take 10 to the -3.8657 power by actually subtracting 10 from 7.8657?

OpenStudy (campbell_st):

because it recognises - as an operation not part of a value... its like -2^2 = -4 on calculators to get it right you need (-2)^2 =4

OpenStudy (anonymous):

but I did put it parentheses 10^(7.8657-10) :(

OpenStudy (campbell_st):

put it down to a programming glitch

OpenStudy (anonymous):

Oh what an idiot I am! 7.8657-10 is not -3.8657, it's -2.1343

OpenStudy (anonymous):

fixing my subtraction gets it right on ti. I will take your advice, assume it's programing and enter exponent as (7.8657-10) . Thanks